OpenStudy (alexanderphantomhive):

Richard works at an ice cream shop. Regular cones get two scoops of ice cream; large cones get three scoops. One hot Saturday, Richard scooped 234 regular cones and 156 large cones. One scoop of ice cream is 3 ounces. A tub of ice cream weighs 10 pounds. How many tubs of ice cream did Richard use to make the cones? a) 936 tubs b) 468 tubs c)175 1/2 tubs d) 17 11/20 tubs

OpenStudy (eliesaab):

Number of pounds = (234 2 + 156 3) 3/16=351/2

OpenStudy (eliesaab):

Number of tubs =(315/2)/10=351/20= 17 11/20

OpenStudy (alexanderphantomhive):

Thank you very much :)